a. The bearing of line A-1 = b. The bearing of line A-2= _N 4E - 30E 250.00! sa i N=1500 200.001 AI Eri 2009
Answer #1

Bearing of line 1-2 = 48°30'

Bearing of line 1-A = 48°30' + 45° = 93°30'

Bearing of line A-1 = 180° +93°30' = 273° 30'

Coordinates of 1 = (1500' N , 1200' E)

Coordinates of 2 = (1665.65' N , 1387.238' E)

Coordinates of A = ( 1487.79'N ,1399.626'E )

Thus length 2-A = 178.29'

Thus 178.29/sin(45°) = 250/sin(A)

Thus A = 82°31'46.49"

Thus bearing of line A-2 = 273°30' + 82°31'46.49" = 356°1'46.49"